# Through the Looking Glass: Mirror Schrödinger Bridges

### v1

@@ Appendix B Analytical Solution for Gaussian Experiment
We follow the proof of (De Bortoli et al., 2021, Proposition 46), which established the corresponding result in the case where the reference process has zero drift. Imitating the proof of (De Bortoli et al., 2021, Proposition 43), we see that the static Schrödinger bridge $\pi^{*}$ exists and is a $2d$-dimensional Gaussian. That the mean equals zero follows from the fact that both marginals have zero mean. The rest of the proof is devoted to determining the covariance matrix $\Sigma$ of $\pi^{*}$.
@@
On the other hand, the mean and variance of the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}$ are computed in (Trajanovski et al., 2023, section II), where it is shown that they are respectively given by
@@
where $C\in\mathbb{R}$ is a nonzero constant independent of $\widetilde{\Sigma}$. As argued in (De Bortoli et al., 2021, proof of Proposition 46), we can assume $S=S^{T}$ is a symmetric matrix, as doing so will only decrease $D_{\rm{KL}}(\widetilde{\pi}\;\|\;\pi^{0})$, so $S$ is diagonalizable. Let $\lambda_{1},\dots,\lambda_{d}$ denote the eigenvalues of $S$, counted with multiplicity. Using the well-known formula for the determinant of a block $2\times 2$ matrix, we find that
@@
Minimizing $D_{\rm{KL}}(\widetilde{\pi}\;\|\;\pi^{0})$ then amounts to take $\lambda_{1}=\cdots=\lambda_{d}=\beta$ in such a way that $f(\beta)$ is minimized. Observe that the equation
@@
Appendix C Implementation Details
In this section we give further details on our experimental setup.
C.1 Gaussian Transport
@@
We use the MLP large network from (De Bortoli et al., 2021) for DSB and DSBM in all Gaussian transport experiments. For our method, we modify this network to take $\sigma$ as an input. The values of $\sigma$ are uniformly sampled from the (inclusive) interval from $1$ to $5$ for training, and at test time we fix $\sigma=1$ for all samples to compare with DSB and DSBM, which do not take $\sigma$ as a network input, but each use $\sigma=1$ via the SDE discretization. We run the same experiment for dimension $d=5$ and $d=20$ (in Figure 6), and $d=50$ (in Figure 1). The number of samples for all experiments is $10{,}000$. We use $20$ timesteps and train for $10{,}000$ inner iterations for each of $20$ outer iterations.
C.2 2D Datasets
@@
We modify the network architecture with positional encoding from (Vaswani et al., 2017), which is used by De Bortoli et al. (2021), to take values of noise $\sigma$ rather than tuples of only ${\mathbf{X}}$ and $t$. The values of $\sigma$ are concatenated to the spatial features before the first MLP block is applied. This modified network is used to parametrize our drift function. We use Adam optimizer with learning rate $10^{-4}$ and momentum $0.9$. We train each example for $10{,}000$ inner iterations per outer iteration of the algorithm. Figure 2 shows the terminal samples obtained for outer iteration $30$ for all example datasets. The noise values $\sigma^{j}$ are sampled uniformly in the range from $1$ to $9$ for training. At test time, a fixed $\sigma$ value is chosen for all sample trajectories. We train with $10{,}000$ samples, which are refreshed each $1{,}000$ iterations. We use $20$ timesteps of size $0.01$ each. All $2$-dimensional experiments run on CPU.
C.3 Image Resampling
For the image dataset experiments, we modify the U-Net architecture used in (De Bortoli et al., 2021; Shi et al., 2023) to take values of noise $\sigma$. Each value $\sigma^{j}$ is expanded to match image size and concatenated to channels of their corresponding sample image $j$ before the input block is applied. For all image experiments we follow the timestep $\gamma$ schedule used in De Bortoli et al. (2021) with $\gamma_{\rm{min}}=10^{-5}$ and $\gamma_{\rm{max}}=0.1$. We use Adam optimizer with learning rate $10^{-4}$ and momentum $0.9$. Experiments with image datasets were run on limited shared GPU resources; lower-resolution image sizes and number of samples in cache were chosen accordingly.
MNIST. For the experiment in Figure 3, we use $10{,}000$ cached images of size $28\times 28$; the batch size is $128$ and the number of timesteps is $30$. The noise values are sampled uniformly in the interval from $1$ to $5$ (inclusive) during training. We train for $5{,}000$ iterations per outer iterations, and cached samples are refreshed every $1{,}000$ inner iterations. The terminal samples shown are for outer iteration $8$.
@@ C.3 Image Resampling
Figure 7: For our generated results (first and seventh columns), we show the five nearest neighbors in the CelebA dataset as measured through the features extracted by ResNet50 (He et al., 2016).
CelebA. In Figures 4 and 8, we use $300$ cached images of size $64\times 64$ and batch size $128$. The cache is refreshed every $100$ inner iterations and we train for $5{,}000$ iterations per outer iterations. The number of timesteps is $50$; the $\sigma$ values are uniformly sampled in the interval from $1$ to $3$. The terminal sample images are shown for outer iteration $15$. The FID score in Figure 4 is computed using $300$ images.
@@
Flowers102. For Figure 5, we use $500$ cached images of size $64\times 64$. The batch size is $128$ and cache is refreshed every $100$ inner iterations. We train for $5{,}000$ inner iterations per outer iteration. Terminal samples are shown for outer iteration $20$. The $\sigma$ values are uniformly sampled in the interval from $1$ to $5$; the number of timesteps is $50$.
Appendix D Additional Experimental Results
(a) Initial Samples
(b) $\sigma=1$
(c) $\sigma=2$
(d) $\sigma=3$
@@ Appendix D Additional Experimental Results
Figure 8: Additional results for the empirical distribution of images in CelebA from which the examples in Figure 4 are obtained.

### v2

@@ Appendix D Analytical Solution to MSB in the Gaussian Case
We follow the proof of Proposition 46 of De Bortoli et al. [2021], which established the corresponding result in the case where the reference process has zero drift. Imitating the proof of Proposition 43 of De Bortoli et al. [2021], we see that the static Schrödinger bridge $\pi^{*}$ exists and is a $2d$-dimensional Gaussian. That the mean equals zero follows from the fact that both marginals have zero mean. The rest of the proof is devoted to determining the covariance matrix $\Sigma$ of $\pi^{*}$.
@@
On the other hand, the mean and variance of the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}$ are computed in section II of Trajanovski et al. [2023], where it is shown that they are respectively given by
@@
where $C\in\mathbb{R}$ is a nonzero constant independent of $\widetilde{\Sigma}$. As argued in the proof of Proposition 46 of De Bortoli et al. [2021], we can assume $S=S^{T}$ is a symmetric matrix, as doing so will only decrease $D_{\rm{KL}}(\widetilde{\pi}\;\|\;\pi^{0})$, so $S$ is diagonalizable. Let $\lambda_{1},\dots,\lambda_{d}$ denote the eigenvalues of $S$, counted with multiplicity. Using the well-known formula for the determinant of a block $2\times 2$ matrix, we find that
@@
Minimizing $D_{\mathrm{KL}}(\widetilde{\pi}\;\|\;\pi^{0})$ then amounts to take $\lambda_{1}=\cdots=\lambda_{d}=\beta$ in such a way that $f(\beta)$ is minimized. Observe that the equation
@@
Appendix E Implementation Details
In this section we give further details on our experimental setup. Akin to Song and Ermon [2020, Technique 5] and De Bortoli et al. [2021, Technique 6], we improve performance of Algorithm 1 by implementing the exponential moving average (EMA) of network parameters. We run our experiments on a NVIDIA GeForce RTX $3090$ GPU with $24$GB of memory.
E.1 Gaussian Transport
@@
We use the MLP large network from [De Bortoli et al., 2021] for DSB and DSBM in all Gaussian transport experiments. For our method, we modify this network to take $\sigma$ as an input. The values of $\sigma$ are uniformly sampled from the (inclusive) interval from $1$ to $5$ for training, and at test time we fix $\sigma=1$ for all samples to compare with DSB and DSBM, which do not take $\sigma$ as a network input, but each use $\sigma=1$ via the SDE discretization. We run the same experiment for dimension $d=5$ and $d=20$ (in Fig. 6), and $d=50$ (in Fig. 2). The number of samples for all experiments is $10{,}000$. We use $20$ timesteps and train for $10{,}000$ inner iterations for each of $20$ outer iterations.
E.2 2D Datasets
@@
We modify the network architecture with positional encoding from [Vaswani et al., 2017], which is used by De Bortoli et al. [2021], to take values of noise $\sigma$ rather than tuples of only ${\mathbf{X}}$ and $t$. The values of $\sigma$ are concatenated to the spatial features before the first MLP block is applied. This modified network is used to parametrize our drift function. We use Adam optimizer with learning rate $10^{-4}$ and momentum $0.9$. We train each example for $10{,}000$ inner iterations per outer iteration of the algorithm. Fig. 3 shows the terminal samples obtained for outer iteration $30$ for all example datasets. The noise values $\sigma^{j}$ are sampled uniformly in the range from $1$ to $9$ for training. At test time, a fixed $\sigma$ value is chosen for all sample trajectories. We train with $10{,}000$ samples, which are refreshed each $1{,}000$ iterations. We use $20$ timesteps of size $0.01$ each. All $2$-dimensional experiments run on CPU.
@@
Figure 7: For our generated results (first and seventh columns), we show the five nearest neighbors in the CelebA dataset as measured through the features extracted by ResNet50 [He et al., 2016].
E.3 Image Resampling
For the image dataset experiments, we modify the U-Net architecture used in [De Bortoli et al., 2021, Shi et al., 2023] to take values of noise $\sigma$. Each value $\sigma^{j}$ is expanded to match image size and concatenated to channels of their corresponding sample image $j$ before the input block is applied. For all image experiments we follow the timestep $\gamma$ schedule used in De Bortoli et al. [2021] with $\gamma_{\rm{min}}=10^{-5}$ and $\gamma_{\rm{max}}=0.1$. We use Adam optimizer with learning rate $10^{-4}$ and momentum $0.9$. Experiments with image datasets were run on limited shared GPU resources; lower-resolution image sizes and number of samples in cache were chosen accordingly.
@@ E.3 Image Resampling
MNIST. For the experiment in Fig. 4, we use $10{,}000$ cached images of size $28\times 28$; the batch size is $128$ and the number of timesteps is $30$. The noise values are sampled uniformly in the interval from $1$ to $5$ (inclusive) during training. We train for $5{,}000$ iterations per outer iterations, and cached samples are refreshed every $1{,}000$ inner iterations. The terminal samples shown are for outer iteration $8$.
CelebA. In Fig. 5 and 12, we use $300$ cached images of size $64\times 64$ and batch size $128$. The cache is refreshed every $100$ inner iterations and we train for $5{,}000$ iterations per outer iterations. The number of timesteps is $50$; the $\sigma$ values are uniformly sampled in the interval from $1$ to $3$. The terminal sample images are shown for outer iteration $15$. The FID score in Fig. 5 is computed using $300$ images.
Appendix F Additional Experimental Results
Figure 8: On the left: Two histograms demonstrating how image samples generated with larger $\sigma$ correspond to less proximal samples relative to the initial image sample. On the right: Two histograms show the inverse ratio between displacement and total path length of sample paths as a metric of path regularity.
F.1 Control Over Sample Proximity
We define proximity of samples using pixel-wise $L_{2}$ norm as our choice of distance metric. In Fig. 8 (left), we demonstrate how larger values of $\sigma$ effectively produce pushforward samples that are farther in this distance metric, compared to samples generated with smaller values of $\sigma$. This experiment expands the results shown in Fig. 3 to the case of resampling from image distributions. In particular, the mean and spread of the histograms in Fig. 8 (right) show that larger values of sigma correspond to higher average distance values relative to the initial sample, as well as greater variation among these distances.
Figure 9: On the left: Three curves, each corresponding to a different $\sigma$ value, showing convergence using Chamfer distance for the same 2D dataset (shown in Fig. 3). On the right: Three curves, each corresponding to a different 2D dataset, showing convergence for a fixed $\sigma$ value.
F.2 Sample Path Regularity
We present empirical results on the regularity of path measures produced by our method. Specifically, in Fig. 8 (right), we give a histogram for the values of a metric defined by taking the ratio of total displacement to total path length for different values of $\sigma$. For a given sample trajectory $\{{\mathbf{X}}_{i}\}_{i=0}^{M-1}$, this metric is explicitly computed by dividing $\|{\mathbf{X}}_{0}-{\mathbf{X}}_{M-1}\|_{2}$ (total displacement) and $\sum_{k}\|{\mathbf{X}}_{k+1}-{\mathbf{X}}_{k}\|_{2}$ (total path length). The greater the value of this metric, the greater the variation in the trajectory; hence, smaller values of this metric are suggestive of greater sample path regularity. We find, as expected, that sample path regularity decreases as $\sigma$ increases.
F.3 Integrity of Initial Distribution
We compute Chamfer distances as a means of measuring the proximity of the pushforward distributions exhibited in Fig. 3 to the corresponding initial distributions. In the mirror case, the pushforward distribution should match the initial distribution, and the Chamfer distance between them should therefore decay as the number of iterations grows. In Fig. 9, we demonstrate how the Chamfer distance decays over outer iterations of our method for the same 2D distribution with different values of $\sigma$ (left), as well as how the Chamfer distance decays for different datasets with fixed $\sigma$ value (right).
Figure 10: On the left: FID scores of pushforward samples versus outer iterations (single run) produced by our method, by DSB, and by DSBM, for a mirror bridge with the MNIST dataset as the marginal distribution. Solid lines correspond to backward models and dashed lines to forward models. On the right: Result of image resampling at outer iteration $20$. For each method and drift direction, the initial samples are displayed on the upper row and the pushforward samples on the lower.
F.4 Comparison to Alternative Methods
We compare our method with DSB and DSBM for image resampling with the MNIST dataset as the initial and final marginal distribution. For this experiment, we use the implementation for DSB and DSBM-IPF available in the code repository for Shi et al. [2023]. We implement our algorithm based on the architecture provided, only modifying the model to take on $\sigma$ as an input parameter for our method. We test all three methods with the same set of training parameters as described in Appendix E.3. We train our model with $\sigma=1$ fixed to match the noise value in the SDE for the other two methods, which do not take $\sigma$ as a model input.
We provide FID scores for each method in Fig. 10. We observe that for DSB and DSBM, the forward and backward models result in pushforward samples of different quality. In particular, sample quality for the forward model is significantly lower than that of the backward. This indicates that neither of these methods converge to the mirror Schrödinger bridge for the given number of iterations, because the drift function for this bridge is necessarily time-symmetric, i.e., the forward and backward drifts must be equal to each other. In contrast, our algorithm provides time-symmetry by construction: a single model is trained and forcibly “symmetrized" at each outer iteration via the drift averaging procedure described in §4.3.
Also in Fig. 10, we present a breakdown of runtime for each method obtained for the same experiment. Our method has significantly lower total runtime and average outer training iteration time. The latter is not surprising, considering that one of the key features of our algorithm is to eliminate training for one of the projection steps taken; recall that we perform the reverse $D_{\mathrm{KL}}$ projection completely analytically. We observe that the average inference time during training, however, is higher with our method. Overall, in this particular experiment, we see that our method makes a trade-off between a small reduction in sample quality for a significant speed-up in training, while also preserving the time-symmetry of the solution.
Table 1: Average runtimes for the experiment in Fig. 10
Ours DSB DSBM-IPF
Total $2.64$ $5.25$ $12.47$ hrs
Outer Iter. $7.94$ $15.7$ $37.41$ min
Inner Iter. $0.059$ $0.055$ $0.209$ s
Inference $2.009$ $1.554$ $1.002$ s
F.5 Additional MNIST Results
@@
Figure 11: Additional results for the empirical distribution of handwritten digits from which the examples in Fig. 4 are obtained.
F.6 Additional CelebA Results
Figure 12: Additional results for the empirical distribution of images in CelebA from which the examples in Fig. 5 are obtained.